OpenStudy (anonymous):

Help please! Find the slope of the straight line that passes through (–2, –4) and (3, –5).

OpenStudy (texaschic101):

slope formula : slope = (y2 - y1) / (x2 - x1) (-2,-4) x1 = -2 and y1 = -4 (3,-5) x2 = 3 and y2 = -5 now we sub slope = (-5 - (-4) / (3 - (-2) slope = (-5 + 4) / (3 + 2) slope = -1/5

OpenStudy (texaschic101):

any questions ?